if the mean deviation about the median of the numbers a,2a,.........,50a is 50,then IaI equals

Grade:10

1 Answers

bharat bajaj IIT Delhi
10 years ago
The median of this list = 25a + 26a / 2 = 25.5a
Mean deviation about the median = |a - 25.5a| + |2a - 25.5a| ....... | 50a - 25.5a| / 50 = 50
50(50)= 25.5a(25) - ( a + 2a... 25a) + (26a + 27a ... 50a) - 25.5a(25)
50(50) = 50(51)a/2 - 25(26)a/2 - 25(26)a/2
50(50) = 25(51)a - 25(26)a
50(2) = 51a - 26a
50(2) = 25a
a=4
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty